OpenStudy (anonymous):

Have you ever worked with variables? Algebra is a branch of mathematics that specifically deals with the idea of using variables to solve problems. What kinds of problems have you already solved using variables? Do you know any real-life applications for variables? Think about the numbers you encounter in the world around you. Do you understand what they really mean? Interactions at grocery stores, gas stations, and online shopping commonly involve numbers. These interactions can be better understood using algebra.

Atsie (atsie):

How funny that you should post this. I'm seriously struggling with variable's right now. I can never seem to find the meaning behind them. Welcome to OpenStudy by the way! :) Your profile pic is gorgeous!

OpenStudy (anonymous):

Absolutely, for example the price of a sandwich is... £2.70. We could call that S. So S=£2.70. If im particularly hungry I might buy two sandwiches which would be 2S, and 2S=2x2.70. Which equals £5.40. Say I bought sandwiches for my whole family, which is 6 people. That would be 6S so 6x2.70. What about a drink? Say my drink cost £1. We could let D=£1. So if my whole family wanted sandwiches and a drink each we could write the cost as... Cost = 6S + 6D Just replace S with 2.70 and D with 1.00 Cost = 6x2.70 + 6x1 Cost = 16.20 + 6 Cost = £22.20